### v1

# Open3DTrack: Towards Open-Vocabulary 3D Multi-Object Tracking

Abstract
3D multi-object tracking plays a critical role in autonomous driving by enabling the real-time monitoring and prediction of multiple objects’ movements. Traditional 3D tracking systems are typically constrained by predefined object categories, limiting their adaptability to novel, unseen objects in dynamic environments. To address this limitation, we introduce open-vocabulary 3D tracking, which extends the scope of 3D tracking to include objects beyond predefined categories. We formulate the problem of open-vocabulary 3D tracking and introduce dataset splits designed to represent various open-vocabulary scenarios. We propose a novel approach that integrates open-vocabulary capabilities into a 3D tracking framework, allowing for generalization to unseen object classes. Our method effectively reduces the performance gap between tracking known and novel objects through strategic adaptation. Experimental results demonstrate the robustness and adaptability of our method in diverse outdoor driving scenarios. To the best of our knowledge, this work is the first to address open-vocabulary 3D tracking, presenting a significant advancement for autonomous systems in real-world settings. Code, trained models, and dataset splits are available at https://github.com/ayesha-ishaq/Open3DTrack.
I INTRODUCTION
3D multi-object tracking involves detecting and continuously tracking multiple objects in the physical space across consecutive frames. This task is essential in autonomous driving as it allows the vehicle to monitor and predict the movements of multiple objects in its environment. The system can make informed decisions for safe navigation, collision avoidance, and trajectory planning by accurately identifying, localizing, and tracking objects over time. This capability is vital for real-time responsiveness and the overall reliability of autonomous systems.
Current 3D tracking systems are structured around benchmarks with labeled datasets, driving the development of methods that excel at tracking predefined categories such as cars, pedestrians, and cyclists [1, 2, 3, 4, 5, 6]. These tracking systems maximize performance within this closed set by detecting and associating known objects across frames. However, these systems are limited by their reliance on a closed set of classes, making them less adaptable to new or unexpected objects in dynamic real-world environments. Although effective within defined benchmarks, their performance diminishes when encountering unknown objects.
To overcome these limitations, open-vocabulary systems are needed to enable the detection and recognition of objects beyond the predefined categories. In 3D tracking, open-vocabulary capabilities are essential for enhancing adaptability and robustness in dynamic environments, where a vehicle may encounter unexpected or rare objects not covered by the training data. By incorporating open-vocabulary approaches, 3D tracking systems can better generalize to new objects, ensuring safer and more reliable autonomous navigation in diverse real-world scenarios.
In this paper, we introduce the task of open-vocabulary 3D tracking, where trajectories of known and unknown object classes are estimated in the real 3D space by linking their positions across consecutive frames. We formulate the open-vocabulary tracking problem, propose various splits to evaluate the performance across known and unknown classes, and introduce a novel approach that bridges the performance gap between tracking known and unseen classes. Our approach leverages the strengths of 2D open-vocabulary methods and existing 3D tracking frameworks to generalize to unseen categories effectively. Through extensive experiments, we demonstrate that our proposed method achieves strong results through our adaptation strategies, providing a robust solution for open-vocabulary 3D tracking. Fig. 1 shows an example output of our solution for the open-vocabulary 3D tracking task. Our method can detect and track previously unseen object categories in three-dimensional space without relying on predefined labels.
Known class track: Pedestrian
Unknown class track: Bus, Bicycle
[…]
Our contributions can be summarised as follows:
  • We introduce the task of open-vocabulary 3D tracking that generalizes to unseen categories.
  • We propose a novel problem formulation and evaluation splits for open-vocabulary tracking.
  • We develop a method that effectively bridges the performance gap between known and unknown class tracking.
  • Our extensive experimental validation shows the high effectiveness of our proposed adaptation.
II Related Work
Our work focuses on 3D tracking and open-vocabulary systems. This section reviews existing approaches in these areas, highlighting key techniques and advancements to provide an understanding of current methodologies.
II-A 3D Tracking
In 3D object tracking, various approaches have been explored, each leveraging different features and methodologies. Traditional approaches often rely on Kalman filters or other probabilistic models, with recent advances shifting towards deep learning-based methods. The tracking-by-detection paradigm remains dominant, particularly in 3D multi-object tracking (MOT). Model-based techniques, such as AB3DMOT [7], extend traditional Kalman filters to 3D for motion estimation, coupling it with the Hungarian method to provide a standard 3D MOT baseline. Another method, EagerMOT [8], combines observations from LiDAR and camera sensors, enhancing the ability to track distant objects and achieve precise trajectory localization. The Immortal Tracker [9] introduces a trajectory prediction mechanism using a Kalman filter to maintain tracks even when objects temporarily disappear from view, significantly reducing identity switches. CenterPoint [3] and CenterTrack [10] combine velocity estimates with point-based detection to track the centers of objects across multiple frames. Poly-MOT [1] introduces a method that adapts a tracking criterion based on object categories, employing category-specific motion models and a two-stage data association strategy. In 3DMOTFormer [2], the authors introduce a geometry-based 3D multi-object tracking framework built on the transformer architecture, using an Edge-Augmented Graph Transformer to perform data association through edge classification. While great advancement has been made in the 3D tracking task, all these methods rely on tracking a known predefined set of object classes.
II-B Open-Vocabulary 3D
Open vocabulary 3D tasks are relatively new and emerging in computer vision, driven by the need to generalize beyond fixed, predefined categories, particularly in dynamic and unstructured environments. Recent advances in 2D open vocabulary tasks [11, 12, 13, 14] have inspired the extension of these concepts to the 3D domain, where the complexity of object shapes, occlusions, and sparse point-cloud data pose unique challenges. Introducing open vocabulary techniques into 3D vision tasks marks a significant shift towards more flexible and scalable models capable of identifying objects across a wide range of categories without requiring exhaustive labeled data. In [15], one of the earlier works, the authors applied transductive zero-shot learning (ZSL) and generalized zero-shot learning (GZSL) to 3D point cloud classification. PointCLIP [16] leverages the Contrastive Vision-Language Pre-training (CLIP) model to achieve open-vocabulary point cloud recognition by aligning CLIP-encoded point clouds with 3D category texts through knowledge transfer from 2D images to 3D. For 3D open-vocabulary tasks, using image features obtained from vision language models and aligning them with point cloud features is common [17], [18]. OpenScene [19] generates dense features for 3D points embedded within the CLIP feature space by using image-per-pixel features from 2D semantic segmentation models [13]. OpenMask3D [20] introduces a zero-shot approach for open-vocabulary 3D instance segmentation by utilizing class-agnostic 3D instance masks and aggregating per-mask features through multi-view fusion of CLIP-based image embeddings. In addition to semantic segmentation, in [21], the authors propose a method that combines a point cloud detector for object localization with cross-modal contrastive learning to connect image, point cloud, and text representations, enabling open-vocabulary 3D detection in indoor scenes. However, open-vocabulary 3D tracking remains largely unexplored, presenting a significant gap in leveraging these emerging techniques for dynamic, real-time object tracking in the physical space.
Figure 2: System overview. Open3DTrack leverages 3D proposals from base classes $C^{base}$ to train the 3D tracker in a class-agnostic manner, enabling it to construct tracks and predict confidence scores. During inference, the system classifies 3D proposals using open-vocabulary categories from both base and novel classes $C^{base}\cup C^{novel}$, utilizing 2D image cues and pretrained vision-language model. This process labels the tracks output by the tracker, effectively generating open-vocabulary 3D tracks.
III Methodology
In this section, we first formulate the problem of 3D multi-object tracking in an open-vocabulary setting. We then describe our approach for adapting an existing tracker to a class-agnostic open-vocabulary tracker. Finally, we outline the dataset splits used for evaluation.
III-A Problem Formulation
Given a sequence of LiDAR frames $\mathbf{P}=\{P_{t}\}_{t=1}^{T}$, where $P_{t}$ represents the point cloud at time $t$, the goal of 3D multi-object tracking is to detect and track multiple objects $\{O_{i}^{t}\}_{i=1}^{N_{t}}$, where $O_{i}^{t}$ denotes the $i$-th object in frame $t$. Each $O_{i}^{t}$ comprises the object class $c$ and the 3D bounding box $b=[x,y,z,w,l,h,\theta]$ where $x,y,z$ represent the coordinates of the center of the bounding box, and $w,l,h$ are the width, length, and height of the bounding box, and $\theta$ is the heading angle, respectively.
In an open-vocabulary setting, we train a tracker $\mathcal{T}$ on $\mathbf{P}^{train}$ and their corresponding annotation $A^{train}$ of objects with semantic categories $\mathcal{C}^{base}$. At test time, we aim to find tracks $G_{T}$ of objects in $\mathbf{P}^{test}$ belonging to $\mathcal{C}^{base}$ as well as a given $\mathcal{C}^{novel}$, where $\mathcal{C}^{base}\cap\mathcal{C}^{novel}=\phi$. In a tracking-by-detection framework, only class labels and scores for $\mathcal{C}^{base}$ object detections are available at train and test time. Our setup requires the tracker $\mathcal{T}$ to track arbitrary object classes with $\mathcal{C}^{novel}$ not appearing during training time.
III-B Open-Vocabulary 3D Multi-Object Tracking
In this section, we present our approach for 3D open-vocabulary tracking. As illustrated in Fig. 2, our proposed system integrates 3D proposals, 2D image cues, and vision-language models to classify and track objects in an open-vocabulary setting, including classes not encountered during training.
3D Tracker
We adapt our 3D open-vocabulary tracking approach from the recent 3D multi-object tracker 3DMOTFormer. [2]. 3DMOTFormer is a tracking-by-detection framework that leverages graph structures to represent relationships between existing tracks and new detections. Features such as position, size, velocity, class label, and confidence score are processed through a graph transformer, and updated based on interactions. Within the decoder, edge-augmented cross-attention models how tracks and detections interact, with edges representing potential matches. To match tracks with detections, affinity scores are predicted which indicate the likelihood of correspondence. The final step computes loss using positive target edges that share the same tracking ID. Ground truth tracking IDs are assigned by calculating 3D IoU between ground truth boxes and 3D detections, followed by Hungarian Matching for one-to-one ID assignment. Loss $\mathcal{L}_{m}$ accumulated over the sequence is used to train the network with back-propagation through time (BPTT). The tracker is trained on a closed set of classes, producing tracks $G_{T}$ of objects from those classes.
While 3DMOTFormer performs well for the closed set, it struggles to track unseen object classes due to its reliance on class-specific information during training. The tracker expects identified 3D boxes to generate edge proposals and then process those through the graph transformer. In a situation where it may receive a new label or an unlabelled object, the model will fail to track it. 3DMOTFormer also uses class-specific statistics within its pipeline, such as maximal class velocities, to compute association edges between tracks and new detections. Such information regarding unseen, unfamiliar objects, that the system may encounter for the first time, is not available in a real-world open-vocabulary setting.
Class Agnostic Tracking
To enable tracking objects irrespective of their class label, we modify the tracker to utilize only object position, dimensions, heading angle and velocity as the initial features. Object detections from a given detector are treated as proposals (class-agnostic) where the class labels and confidence scores are dropped and not utilized.
To compute potential association edges, we replace class-specific distance thresholds with a fixed threshold, achieving a balance between maintaining true edges and avoiding unnecessary associations.
We change the GT track ID assignment by computing 3D IoU between ground truth boxes $A^{train}$ and 3D detections belonging to $\mathcal{C}^{base}$ in a class-agnostic manner.
Figure 3: Open-vocabulary dataset splits. Our proposed dataset splits for known and unknown categories are based on statistics from the nuScenes[22] dataset, which are shown here. The first split considers the least occurring objects as novel. In the second split, we analyze the motion patterns of objects in urban settings, motivating our urban vs highway split. Novel classes in this split exhibited the highest percentage of movement. In the final split, we use the average points per box to ensure a similarly diverse object distribution in the novel split in terms of object volume.
Confidence Score Prediction
As discussed earlier, we drop class labels and confidence scores from the detection features to ensure the tracker is independent of any information specific to object classes. As a result, the initial output from the system suffers from a lack of representation of the objectness score of each box. To overcome this issue, we use the confidence scores available for $\mathcal{C}^{base}$ and train a regression head at the end of the tracker. The regression head is composed of a feed-forward network that takes as input processed detection features $h^{(Ld)}_{D,j}$ from the graph transformer and predicts a single confidence score. We employ a straightforward Mean Squared Error loss between predicted and true confidence scores to learn the scores for the detection boxes. The combined loss is $\mathcal{L}=\mathcal{L}_{m}+\lambda_{c}\mathcal{L}_{c}$, where $\mathcal{L}_{c}$ is the loss for confidence score, weighted by $\lambda_{c}$.
2D Driven Open Vocabulary Labels
While we train the tracker by removing all class-specific dependencies, we require a strong open-vocabulary classification for all object tracks, known and unknown, at test time. We leverage vision-language models to match regions in the images with text prompts, enabling it to detect and classify objects for which it has not been explicitly trained. 2D detections on multi-view images from the vehicle are obtained from an open-vocabulary 2D detector, which is prompted with all possible labels—both $\mathcal{C}^{base}$ and $\mathcal{C}^{novel}$. The system allocates class labels to the 3D proposals by first projecting the 3D boxes onto multi-view images. The labels are then derived from the highest IoU overlap with 2D detections on the image plane. Projections of the 3D detections that do not overlap with a 2D open-vocabulary detection are labeled ’unknown’. We explicitly identify these boxes at the end of the tracking process.
Track Consistency Scoring
At the output of the tracker, we apply a track smoothing module that addresses the issue of unknown boxes that do not match any 2D detections from the images, as well as improve label consistency throughout tracks to account for inaccuracies in the open-vocabulary detector outputs and errors accrued during projection from 3D to 2D. To calculate the most accurate class for each track, we first compute the weight of an object based on its detected bounding box in an image, with objects at a greater distance assigned a smaller weight. The depth is derived from the size of the bounding box, the object’s vertical position in the image for perspective correction, and the aspect ratio of the bounding box to ensure uniformity for all classes.
$B_{size}=\frac{B_{width}\times B_{height}}{I_{width}\times I_{height}}$
$depth=\frac{1}{B_{size}}(1-\alpha_{p}\left(\frac{y_{center}}{I_{height}}\right))$
$depth=\begin{cases}depth\times\left(\frac{1}{aspect\_ratio}\right)&\text{if }aspect\_ratio>\beta_{ar}\\ depth&\text{otherwise}\end{cases}$
where, $B$ refers to 2D detected box, $I$ refers to image, $y_{center}$ points to the vertical center of the box, $\alpha_{p}$ is the perspective correction factor, while $\beta_{ar}$ is the aspect ratio threshold and
$aspect\_ratio=\frac{B_{width}}{B_{height}}$
We then compute the weight as
$w_{dist}=e^{\frac{-depth}{\lambda_{s}}}$
where $\lambda_{s}$ is a depth scaling factor. The predicted confidence score of each box is multiplied with $w_{dist}$, yielding higher confidence for boxes that appear closer in images. The smoothing module then uses the modified confidence score and finds the average modified confidence for each class in the track, along with their occurrence, selecting the class with the highest combination. Any track that is composed entirely of unknowns is dropped.
III-C Open-Vocabulary Tracking Splits
Our method is evaluated in outdoor environments using custom data splits based on common object types and situations. Fig. 3 shows various statistics from nuScenes [22] and subsequent splits. One split follows the common open-vocabulary approach, treating rarer classes as novel. Another split differentiates between highway and urban objects based on the greater activity and mobility of some classes within urban areas. A final split focuses on objects with high variability in shape and size, represented by the average LiDAR points per object. These splits provide a robust framework for evaluating open-vocabulary tracking in realistic and diverse outdoor settings.
IV Experiments
[…]
Overall Bicycle Bus Car Motorcycle Pedestrian Trailer Truck
AMOTA AMOTP AMOTA AMOTP AMOTA AMOTP AMOTA AMOTP AMOTA AMOTP AMOTA AMOTP AMOTA AMOTP AMOTA AMOTP
Split 1 - Rare
Baseline
(w/o unknowns)
0.415 0.868 0.323 1.051 0.403 1.030 0.656 0.432 0.194 1.336 0.717 0.399 0.395 1.069 0.218 0.758
Baseline
(w/ unknowns)
0.483 0.785 0.424 0.853 0.423 0.952 0.678 0.391 0.288 1.293 0.746 0.350 0.469 0.975 0.349 0.684
Open3DTrack
0.578 0.783 0.445 0.988 0.612 0.918 0.779 0.390 0.469 1.104 0.752 0.403 0.477 1.037 0.511 0.638
Split 2 - Urban
Baseline
(w/ unknowns)
0.469 0.740 0.301 0.982 0.481 0.907 0.657 0.392 0.537 0.549 0.482 0.634 0.455 0.985 0.367 0.729
Open3DTrack
0.590 0.677 0.400 0.894 0.683 0.820 0.788 0.387 0.702 0.422 0.548 0.599 0.488 0.981 0.522 0.635
Split 3 - Diverse
Baseline
(w/ unknowns)
0.438 0.813 0.389 0.493 0.640 0.499 0.627 0.402 0.275 1.311 0.564 0.590 0.424 0.993 0.144 1.406
Open3DTrack
0.536 0.804 0.524 0.581 0.770 0.543 0.708 0.395 0.438 1.143 0.564 0.648 0.470 1.036 0.276 1.281
Upper Bound
0.710 0.521 0.545 0.429 0.853 0.527 0.838 0.382 0.723 0.459 0.812 0.335 0.509 0.936 0.690 0.576
IV-A Datasets
We use the nuScenes [22] dataset for training and evaluation of our open-vocabulary tracking system. NuScenes features data collected from several sensors, such as 360-degree cameras, a 32-beam LiDAR, and RADARs annotated at 2Hz. It includes 1000 scenes, each lasting 20 seconds, and is split into 700 scenes for training, 150 for validation, and 150 for testing. It offers seven unique classes for the tracking task.
Evaluation Metrics
We evaluate our results using the nuScenes [22] tracking benchmark with AMOTA and AMOTP as the primary metrics [23]. AMOTA averages MOTA across recall levels, accounting for false positives, missed targets, and identity switches, with lower impact at recall levels below 10%. AMOTP measures the average alignment error between predicted and annotated bounding boxes, reflecting localization accuracy across all frames.
IV-B Implementation Details
We follow 3DMOTFormer’s default settings, training for 12 epochs with AdamW, a batch size of 8, and a learning rate of 0.001 with 0.01 weight decay. The regression loss weight $\lambda_{c}$ is set to 0.5. In the association graph, the distance threshold for edge truncation is 3 meters. For open-vocabulary 2D detection, we use the YOLOv8L-Worldv2 model [24] with a confidence threshold of 0.01. In the Track Consistency Scoring module, perspective correction $\alpha_{p}$ is 0.2, aspect ratio threshold $\beta_{ar}$ is 2.5, and depth scale $\lambda_{s}$ is 250.
IV-C Benchmark Results
Figure 4: Qualitative comparison. We show the output of our method and the baseline on challenging examples from nuScenes. The top two rows show an occluded novel class object, bus, which is tracked successfully by ours, while it changes IDs in the baseline output. The bottom two rows show an example of poor lighting and blurred novel object, pedestrian, tracked by ours but generating a fragmented track by the baseline. Additional results are presented in the accompanying video.
Table I shows results on the nuScenes [22] validation set using 3D proposals from CentrePoint [3] detector. We compare our open-vocabulary method on the various proposed splits with the baseline. For reference, we show the original closed-set results of 3DMOTFormer as an upper bound. The baseline method is a simple adaptation of 3DMOTFormer, where we replace class labels with labels obtained from 2D open-vocabulary detections and make use of confidence scores from the 2D detections as well. Baseline without unknowns drops any 3D proposals unmatched to 2D detections, while baseline with unknowns forms tracks with them and assigns them labels of the most common class within the track. For the first split, we show results for baseline with and without unmatched detections. With our presented adaptations, we achieve an overall AMOTA of 0.567 in the first split, with strong performance scores for the novel classes, improving significantly on top of the baseline. In each of the last two splits, we achieve a similar improvement over the baseline, with an overall AMOTA of 0.59 in split 2 and 0.536 in split 3. Particularly in split 3, we notice the performance being sub-par for novel class trucks; this is due to the vast variations in appearance, size, and type of vehicles for class trucks within the dataset, making this category a considerably difficult unknown class. We show some challenging novel class examples in Fig. 4.
TABLE II: Results on the nuScenes [22] validation set for detector generalization for open-vocabulary 3D tracking. Results are shown for base and novel class groups from split 3 III-C, with the tracker trained on detections from the source method and tested on 3D detections from the target method.
Base Novel Overall
Source Target AMOTA AMOTP AMOTA AMOTP AMOTA
Centerpoint Centerpoint 0.630 0.617 0.509 1.003 0.578
MEGVII MEGVII 0.586 0.7055 0.477 1.1207 0.539
BEVFusion BEVFusion 0.662 0.5535 0.410 1.054 0.554
Centerpoint MEGVII 0.579 0.7106 0.496 1.093 0.543
Centerpoint BEVFusion 0.616 0.556 0.458 0.982 0.548
We evaluate our method’s generalization across different detectors, CenterPoint[3], MEGVII [25] and BEVFusion [26], as shown in Table II. Our approach delivers consistent performance across various 3D proposal sources. The second part of the table shows generalization results when trained on proposals from one detector and tested on another. The best results are from the standard detector, likely due to alignment between train and test bounding boxes. Interestingly, while BEVFusion gives the highest base class performance, it results in the lowest novel class performance, possibly due to focusing on high-quality base class detections and less noise for regulation effect.
TABLE III: Ablation study on the effect of each component of our proposed method.
Component Base Novel
AMOTA AMOTP AMOTA AMOTP
Baseline 0.504 0.6638 0.482 0.841
+ Class-Agnostic GT assignment 0.536 0.6215 0.437 0.804
+ Confidence Score Prediction 0.618 0.617 0.507 0.7987
+ Track Consistency Scoring 0.625 0.6063 0.544 0.771
V Ablation
In this section, we detail the effects of various components of our proposed system. All experiments reported are on the nuScenes validation set using CenterPoint detections and specifically on split 2 from section III-C. Table III displays the improvements with each component.
Baseline: As detailed in section IV-C, baseline for our system is simply 3DMOTFormer trained on labels obtained from 2D open-vocabulary detections. All other class-specific details were removed. A simple track smoothing method is applied to the results.
Class Agnostic GT Assignment: To build on top of our baseline we assign GT tracking IDs to 3D proposals using merely 3D IoU, omitting class labels. This improves results for the base classes at test time while results drop slightly for novel classes. This phenomenon occurs because class agnostic GT assignment improves target accuracy as compared to noisy open-vocabulary base class labels, leading to better learning on known classes, while reducing the regularization and dropping performance on unknown classes. However, the overall improvement in performance is significant.
Confidence Score Prediction: As discussed in section III-B, the confidence scores from the 2D detections did not capture the true objectness score of the 3D proposals resulting in lower performance. There was also no estimation of confidence scores for unmatched proposals. To overcome these issue we added the confidence prediction head. We see a significant boost from this module in both the base and novel classes, 8.2%P and 7%P respectively
Track Consistency Scoring: Simple track smoothing is necessary to assign labels to unknown detections, however, we found that formalizing a mechanism to score the identified detections and use these scores to attain consistency in track labels considerably improved performance for all categories. The overall AMOTA increased by 2%P upon the addition of this module.
VI Conclusion
In conclusion, we introduce a novel approach for open-vocabulary 3D multi-object tracking, addressing the limitations of closed-set methods. By leveraging 2D open-vocabulary detections, and applying strategic adaptations to a 3D tracking framework such as class-agnostic tracking and confidence score prediction, we demonstrate improvements in tracking novel object classes. Our experiments highlight the robustness of our method. These contributions represent a meaningful step toward more flexible 3D tracking solutions for autonomous systems, enhancing their ability to generalize in dynamic environments.